1/1

Lord George Gordon Byron ,

My Soul is Dark

My soul is dark - Oh! quickly string The harp I yet can brook to hear; And let thy gentle fingers fling Its melting murmurs o’er mine ear. If in this heart a hope be dear, That sound shall charm it forth again: If in these eyes there lurk a tear, ’Twill flow, and cease to burn my brain.

But bid the strain be wild and deep, Nor let thy notes of joy be first: I tell thee, minstrel, I must weep, Or else this heavy heart will burst; For it hath been by sorrow nursed, And ached in sleepless silence, long; And now ’tis doomed to know the worst, And break at once - or yield to song.

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1110-2240 (2015-01-03 22:28:23)

Tiun ĉi anglan poemon mi, Manfred Retzlaff, trovis en la retejo http: // classiclit. about. com/ library bl-etexts/ lbyron/ bl-lbyron-mysoul. htm .

http://www.poezio.net/version?poem-id=1110&version-id=2240